A car moving south speeds up from 10 m/s to 40 m/s in 15 seconds. What is the car’s acceleration?

Grade:12

1 Answers

Abesh Roy
6 years ago
We know that ,Acceleration = change in speed per unit timeInitial Speed = 10 m/secFinal Speed = 40 m/secChange in speed= (40-10)m/sec = 30m/secTime taken to accelerate = 15 secsSo, acceleration = change in speed / time =30 m/sec / 15 sec =30/15 m/sec2 =2 m / sec2